(17 points) Match the third order linear equations with their fundamental solution sets. 1. y - Y - +y=0 2. y - 5y + 6
Answer #1

1. y - y - ylty = 0 (m3-m²-m -m²-m+1) y = m3-m²- dian m - m - m t -1 - m=+1 J o ) 1 - o wot are (- رارا t te -t te E solun sey mtmo 3 M( n + 1) =0 m=o, m²=t set, I, cost, sint 9 ③ y by tyl by a m3-6m²tm-6 - m²(m-6)+1(m-6)=0 (m²+1)=0 (n-6)=0 ti m =